OpenStudy (anonymous):

For each of the following complex numbers give the, absolute value; the opposite; the conjugate: 1) 3-4i 2) 8i

OpenStudy (anonymous):

the absolute value is the length of the complex number when we think of it as a vector: eg 2 + 6i  we use Pythagoras' theorem \[|2 + 6i| = \sqrt{2^2 + 6^2} = \sqrt{40}\] the conjugate is the same complex number z, but we change the sign of the imaginary part, common notation is z* is the conjugate of z for z = 2 + 6i z* = 2 -6i when you say opposite do you inverse?